Please help me solve the quadratic formula
9/8-2=5/x^2
Thank you for your time

Answer by bobdabuildr(27)

(9/8)-2=5/(x^2)
(-7/8)=5(x^2) Cross muliply
7x^2=-40 Divide both sides by 7
x^2=(-40/7)find the sqrt of both sides
x= i*sqrt(40/7)
Unless you ment to write the problem as 9/(8-2)=5/(x^2) and then you will have a different answer, but the same basic steps.